OpenStudy (anonymous):

Multiply and simplify. 2/3 x(1/3+3 1/4)

OpenStudy (studygurl14):

Hello @baconbaby1334 ! Okay, so your first step would be to convert \(\Large 3\frac{1}{4}\) into an improper fraction. Can you do that?

OpenStudy (anonymous):

13/4?

OpenStudy (michele_laino):

ok!

OpenStudy (studygurl14):

Correct. So now you have \(\Large \frac{2}{3}(\frac{1}{3}+\frac{13}{4})\) Now, there is two ways to do this. You can add the stuff in the parenthesis first and then multiply that by 2/3, or you can distribute the 2/3 to the stuff in the parenthesis, and then add. Which way do you want to do?

OpenStudy (anonymous):

parenthesis

OpenStudy (shaik0124):

2/3(43/12) 43/18 is answer

OpenStudy (studygurl14):

Okay, so to add \(\frac{1}{3}\) and \(\frac{13}{4}\) you need them to havet he same denominator. do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

Okay, so what do you get when you add them together?